# Intentional Gesture: Deliver Your Intentions with Gestures for Speech

### v1

@@ 6.2 Evaluation of Intentional Gesture Tokenizer
Effect of Intentional Semantic Supervision. We design three supervision strategies using H-AuMoCLIP features: (a) CNN-layer supervision for low-level spatial detail, (b) transformer-layer with mean pooling for high-level sequence supervision, and (c) transformer-layer with frame-level temporal supervision. As shown in Table 3 (Left), integrating intentional supervision into motion token learning significantly improves generation quality—both in FGD and diversity—while maintaining strong reconstruction fidelity (L1, rFGD). Compared to CNN-only supervision, transformer-based supervision better captures communicative semantics, validating the benefit of higher-level grounding. Notably, frame-level temporal supervision achieves the best overall FGD and diversity, demonstrating that temporal alignment of intention and motion is key to learning expressive gesture representations.
6.3 Evaluation of Gesture Motion Generation
Design Analysis. Replacing spatial-temporal attentions in GestureLSM with self-cross attention and DiT [24] based architecture improves the generation performance on all-speaker setting. We further investigate how intention representations impact gesture generation. As a baseline, we condition the generator on sentence-level BERT embeddings, which provides only marginal improvement (Table 3, Right). Replacing these with intention embeddings derived from H-AuMoCLIP—aligned to audio and semantic context—yields substantial gains in FGD and diversity.
Sensitivity Analysis. We evaluate the robustness of our model under noisy or mismatched intention inputs. Using LLM-rewritten paraphrases of intention descriptions yields negligible differences in FGD and diversity, suggesting that the model is not sensitive to surface-level textual variation. When using temporally unmatched or missing intention text, FGD degrades slightly (by 0.025–0.03), and synchronization weakens marginally, but the model remains stable overall. We attribute this to our alignment mechanism: we extract intention-aligned segments based on transcript word timestamps and apply a temporal overlap threshold (0.5 or 0.7) to guide supervision.
7 Conclusion
We present Intentional Gesture, a novel framework that formulates gesture generation as an intention reasoning task grounded in high-level communicative functions. We first curate the InG dataset by augmenting BEAT-2 with structured annotations of latent intentions—such as emphasis, comparison, and affirmation—extracted via large language models. Then, we introduce the Intentional Gesture Tokenizer, which discretizes global motion while embedding intention semantics into the representation space through targeted supervision. By bridging speech semantics and motion behaviors, our method produces gestures that are both temporally synchronized and semantically meaningful. Experiments on BEAT-2 demonstrate strong improvements in gesture quality, offering a controllable and modular foundation for expressive gesture generation in digital humans and embodied AI.

### v2

@@ Appendix E Additional Experiments
Long Sequence Generation Quality.
In the main paper, the experiment setting were conducted to generate sequences for the whole testing sequence. Specifically, we follow the existing works Liu et al. (2023; 2025a); Chen et al. (2024b) to utilize a sliding window for long sequence generation (with an average of 65.66 seconds). Each time, we provide the previous 2.13 seconds (a sequence length of 16 for neural representation) generated from the previous generated segment as the condition for the current time segment. Naturally, this setting is easy to encounter the error propagation issue (if the sequence from the previous generation present low quality, this error will be propagated to the current time segment). To understand this effect, we further design the new setting that replicate the inference setting of the same inference audio length as that utilized during training (8.633 seconds). We present the comparison setting between EMAGE, GestureLSM and Intentional-Gesture for single speaker setting in Tab.13. On long sequences, our model achieves the best performance (FGD = 0.379, BC = 0.690, Div. = 11.00). Under short‐sequence inference, our FGD further improves by 0.133 (to 0.246), closely matching the improvements of 0.140 and 0.107 seen for EMAGE and GestureLSM, respectively—indicating a consistent FGD gap of  0.12 across models. Note that BC is not reported (–) for 8.633 s segments, due to the tricky implementation to select the precise audio segments from full ground-truth sequences with the generation segments. These results underscore the impact of error accumulation in sliding‐window co‐speech gesture generation and motivate future work on mitigating segment‐wise propagation.
Table 12: Ablation on Speaker Diversity. Increasing speaker diversity consistently boosts retrieval for both seen (Known) and unseen (Unknown) speakers, indicating better generalization.
Known Unknown
[…]
1 20.63 40.34 50.67 1.03 1.95 2.56
2 29.41 57.63 60.61 1.44 2.31 2.78
3 31.37 60.42 63.39 1.67 2.49 2.92
4 33.52 63.52 66.87 1.87 2.64 3.01
Table 13: Comparison of long‐sequence (full test sequences) vs. short‐sequence (8.633 s) inference on the single‐speaker setting.
Long-seq Generation Short-seq Generation
[…]
GT $0.703$ $11.97$ $0.703$ $11.97$
Single-speaker
EMAGE Liu et al. (2023) $0.570$ $0.793$ $11.41$ $0.430$ - $9.57$
GestureLSM Liu et al. (2025a) $0.408$ $0.714$ $13.24$ $0.301$ - 12.12
Ours 0.379 0.690 11.00 0.246 - $10.21$
Quantizer Comparisons Analysis
To isolate the influence of architecture on tokenizer performance, we standardized all encoder–decoder backbones to our CNN+Transformer design, which we found consistently outperforms alternatives across various quantizers. Specifically:
(1) EMAGE Liu et al. (2023) originally uses separate VQ quantizers for upper body, lower body, and hands. We replace its CNN encoders with our ResNet-style CNN blocks and normalize codebook embeddings rather than using raw outputs. We keep the original codebook size and dimensionality to demonstrate how reducing dimension and increasing code count affects performance.
(2) For RAG-Gesture Mughal et al. (2025), we re-implement their encoder and decoder based on Latent Motion Diffusion from MotionLCM codebase Dai et al. (2024). The comparions indicates for the continuous representation, it is hard to present the motion latent with a compressed latent mean prediction from VAE encoder to ensure it is synchronized with the audio for generation.
(3) For ProbTalk Liu et al. (2024b), we maintain their design of product quantization while improve the encoder and decoder with our design. This comparison indicates the product quantization, while present a latent codebook split, unlike our codebook design of separate latent motion representation, presents an inferior performance.
(4) For GestureLSM Liu et al. (2025a), we maintain the design of 6 layers of codebooks for each body region (upper, lower and hands), which leads to 18 codebook in total. While this multi-codebook approach achieves competitive reconstruction, its reliance on separate decoders for sequential region generation reduces efficiency and harms overall motion quality.
Figure 10: User Study Screenshot
Appendix F User Study Details
For user study, we recruited 20 participants with good English proficiency. To conduct the user study, we randomly select videos from GestureLSM Liu et al. (2025a), EMAGE Liu et al. (2023), CAMN Liu et al. (2022b) and ours. Each user works on 8 videos. The users are not informed of the source of the video for fair evaluations. A visualization of the user study is shown in Fig 10.
Appendix G Metric Details
Fréchet Gesture Distance (FGD).
We adopt Fréchet Gesture Distance Yoon et al. (2020) to quantify the distributional similarity between real and generated gestures. Inspired by FID in image generation, FGD compares mean and covariance statistics of latent features extracted from a pretrained network:
$\text{FGD}=\left\|\mu_{r}-\mu_{g}\right\|^{2}+\text{Tr}\left(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{1/2}\right),$ (5)
where $(\mu_{r},\Sigma_{r})$ and $(\mu_{g},\Sigma_{g})$ are the empirical means and covariances of real and generated gesture embeddings, respectively. Lower FGD indicates better realism and distributional alignment.
L1 Diversity (Div.).
To assess sample-level variation, we compute L1 Diversity Li et al. (2021a), defined as the average pairwise L1 distance across $N$ generated sequences:
$\text{L1 Diversity}=\frac{1}{2N(N-1)}\sum_{t=1}^{N}\sum_{j=1}^{N}\left\|p_{t}^{i}-\hat{p}_{t}^{j}\right\|_{1},$ (6)
where $p_{t}^{i}$ and $\hat{p}_{t}^{j}$ denote the joint positions at frame $t$ for the $i$-th and $j$-th sequences. To focus on local articulation, global translation is removed before computing distances.
Beat Constancy (BC).
Beat Constancy Li et al. (2021b) measures rhythmic alignment between gesture dynamics and speech. Motion beats are detected as local minima in upper body joint velocity, while speech onsets define audio beats. BC is computed as:
$\text{BC}=\frac{1}{|g|}\sum_{b_{g}\in g}\exp\left(-\frac{\min_{b_{a}\in a}\|b_{g}-b_{a}\|^{2}}{2\sigma^{2}}\right),$ (7)
where $g$ and $a$ are the sets of gesture and audio beats, respectively. BC closer to ground-truth implies stronger gesture-speech synchronization.
Appendix H Ethical Statement
While our work is centered on generating human motion videos, it raises ethical concerns due to its potential misuse for photorealistic human motion retargeting. We emphasize the importance of responsible use and recommend implementing practices such as watermarking and deepfake detection to mitigate the risks involving deepfake videos and animated representations.
Appendix I Reproducibility statement
We have provided the code of algorithmic annotation for the motion pattern analysis in the supplementary material together with the code for the whole system.
Appendix J The Use of Large Language Models
We utilize Large Langauge Models for the dataset annotation and paper polishing.
Appendix K Limitations
While our framework demonstrates strong performance across alignment, tokenization, and gesture generation, several limitations remain.
First, our method relies on pre-annotated sentence-level intention descriptions to guide semantic learning. This setup assumes that such annotations are either available or can be reliably extracted, which may not hold in less curated or low-resource scenarios. Future work could explore unsupervised or weakly supervised intention discovery to broaden applicability.
Second, while the multi-codebook tokenizer introduces structure into the latent space, it does not guarantee complete disentanglement between semantic and rhythmic dimensions. Investigating more principled inductive biases or factorized token learning may improve interpretability and controllability.
Third, as shown in Sec.E, we discover that existing methods present error propagation issues for long-sequence generation settings. We would like to highlight this issue and hope future works can propose solutions for this fundamental issue for the co-speech gesture generation domain.
Fourth, in this work, while the motion description annotation, gesture-behavior function mapping are intermediate outputs during the annotation procedure, they are not input as variables for the motion control but only intention annotations were utilized. We build this simple baseline because during inference procedure, we are not able to obtain these motion relationed analysis. However, we argue that the values of these annotations should not be ignore and hope future works can further explore the use cases of these annotations as well for motion control and inspire the analysis of the relationships between gesture motion patterns and linguistic cues from speech context.
Finally, although our hierarchical alignment improves generalization across speakers, domain shifts—such as significant accent variation, disfluency, or cultural gesture norms—remain challenging. Incorporating domain adaptation techniques or cross-cultural gesture modeling could enhance robustness in real-world deployments.